### v1

@@ 3.3 Flux magnification
Since the luminosity distance is affected by perturbations (Bonvin et al. 2006; Challinor & Lewis 2011), the measured fluxes will also be affected by relativistic effects such as magnification and Doppler corrections. While the observed fluxes in FS include observational effects such as dust extinction, magnification due to lensing and Doppler effects are not included. We corrected the fluxes to account for these effects using the relation
@@
$\mu_{\rm L}=\frac{1}{(1-\kappa)^{2}-\gamma_{1}^{2}-\gamma_{2}^{2}}\,$ (29)
@@ 3.4 Flux reference to perform split in bright and faint
In order to ensure that bright and faint populations have the same redshift evolution across all bins, we do not use a single flux cut. In fact, flux is inversely proportional to the square of the luminosity distance, meaning that a fixed flux cut would lead to more bright objects at low redshift and more faint objects at high redshift. To avoid this asymmetry in the redshift distribution, we perform the split in $20$ sub-bins, following a similar procedure to the one described in Bonvin et al. (2023).
@@ 3.5 Estimation of biases
We describe here the method we use to estimate the galaxy bias and local count slope for the full Euclid spectroscopic sample, as well as for the bright and faint populations. In our analysis, we neglect the impact of evolution bias.
@@ 3.5.1 Galaxy bias
The galaxy biases for the full and the bright/faint samples are estimated using the method described in Lepori et al. (2023) and Bonvin et al. (2023). We construct a map of the galaxy number counts using HEALPix (Górski et al. 2005) and we extract its angular power spectrum; and we fit it to a theory prediction of the power spectrum computed with the code class (Di Dio et al. 2013, 2014b; Lesgourgues 2011; Blas et al. 2011), using HMCODE to account for nonlinear effects (Mead et al. 2016). The angular power spectrum of the map is estimated with the library NaMaster (Alonso et al. 2019), which corrects for the effect of the mask. The HEALPix map is constructed from the comoving position of the objects, and therefore its angular power spectrum only includes density contributions. Thus, we fit the galaxy bias in real space, so include only the density term in our theory model. In the fit of the simulated spectrum to the model, we fix the cosmology so that the only free parameter is the galaxy bias.
@@ 3.5.2 Local count slope
and we have introduced the LF of bright and faint galaxies in the redshift bin $\bar{z}$, that is $\bar{N}_{\mathrm{B}}\equiv N_{\mathrm{B}}(\bar{z},\,\bar{L}_{\mathrm{H}\alpha}\geq{L}_{\mathrm{H}\alpha,\,\mathrm{lim}},\,\bar{L}_{Y_{\scriptscriptstyle\rm E}}\geq{L}_{Y_{\scriptscriptstyle\rm E},\,\mathrm{lim}})$ and $\bar{N}_{\mathrm{F}}\equiv N_{\mathrm{F}}(z,\,\bar{L}_{\mathrm{H}\alpha}\geq{L}_{\mathrm{H}\alpha,\,\mathrm{lim}},\,\bar{L}_{Y_{\scriptscriptstyle\rm E}}<{L}_{Y_{\scriptscriptstyle\rm E},\,\mathrm{lim}})$, respectively. The luminosities $\bar{L}_{X}$ are the luminosities corresponding to the flux $F_{X}$ in the background FLRW cosmology.
@@
In practice, we estimate the slopes $s_{\mathrm{B/F},\mathrm{H}\alpha}$ by fixing the flux cuts in the $Y_{\scriptscriptstyle\rm E}$ band and performing the same cuts on a version of the mock that is deeper in $\mathrm{H}\alpha$ flux. The slope $s_{\mathrm{B},Y_{\scriptscriptstyle\rm E}}$ is estimated after applying the selection in the $\mathrm{H}\alpha$ flux, from the cumulative LF of the bright population at the flux cut in the $Y_{\scriptscriptstyle\rm E}$ band. Since we perform the split between bright and faint in sub-bins, the $Y_{\scriptscriptstyle\rm E}$-band flux cut is redshift dependent. We estimate the slope of the cumulative LF in sub-bins and estimate an effective value by taking an average weighted by the number of objects in each sub-bin.
@@ 3.6 Measurements of the dipole and its covariance
We use the Landy–Szalay (LS) estimator (Landy & Szalay 1993) to measure the 3D cross-correlation between bright and faint objects within each redshift bin,
@@
The LS estimator has been shown to have the lowest possible variance and bias among all possible estimators based on pair counts, if the correlation is small ($\xi\ll 1$) and if the number densities of the random catalogues are sufficiently high (Landy & Szalay 1993; Kerscher et al. 2000; Keihänen et al. 2019). In order to get unbiased results, we hence create random catalogues that have ten times as many objects as their corresponding data catalogues.
In practice, we create the Poisson-sampled random catalogues from the density distribution of the data catalogues using the inversion method, as described in Schulz (2023). The LS estimator is computed using a modified version of the publicly available code CUTE (Alonso 2012). The modifications, which are the same as in Breton et al. (2019), extend the capabilities of CUTE to allow the computation of odd multipoles in the 2-point correlation function. We count the pairs for separations $2\,h^{-1}\,\mathrm{Mpc}\leq d\leq 30\,h^{-1}\,\mathrm{Mpc}$ and orientations $-1\leq\mu\leq 1$, binning into $d$-bins of width $\Delta d=2\,h^{-1}\,\mathrm{Mpc}$ and $\mu$-bins of width $\Delta\mu=0.004$.
We do not estimate the correlation for smaller separations because such separations are not well sampled by our catalogues. The average separation between a bright object and the closest faint object can be derived from the cross-correlation function monopole, as it represents the average overdensity of faint objects in the neighbourhood of a bright object. From the measurements of the real-space monopoles on the light cone at small separations we find that there are typically no faint objects separated by $d\lesssim 1\,h^{-1}\,\mathrm{Mpc}$ from the average bright object, and at $\bar{z}=1.0$, the average separation between a bright and the closest faint object is $d=3.6\,h^{-1}\,\mathrm{Mpc}$. The number of pairs with separation below this value is generally small, leading to very large Poisson errors on the measurement at those separations. Additionally, at very small separations, the LS estimator can in principle become undefined in certain ($d,\mu$)-bins if there are no pairs between the two random catalogues. For these reasons, we exclude $d<2\,\,h^{-1}\,\mathrm{Mpc}$ from the analysis.
@@
In practice, the integral is numerically computed as a sum over the $\mu$-bins. We thus compute the dipole of the LS estimator as
@@
The covariance of the measurement is estimated with the jackknife (JK) method (Norberg et al. 2009),
@@
where the survey volume at redshift $\bar{z}$ is divided into $N=50$ sub-volumes with approximately equal surface area and shape by running a kmeans [^3] https://github.com/esheldon/kmeans_radec/ algorithm (Kwan et al. 2017; Suchyta et al. 2016) on one of the corresponding random catalogues. The same split is applied to the data catalogues and the LS dipole is measured $N$ times, each time removing a different sub-volume $k$ from the survey. In Eq. (36), $\xi^{k}_{1}(\bar{z},d_{i})$ is the dipole measured within the survey volume at redshift $\bar{z}$ from which the $k$-th sub-volume has been removed, and $\bar{\xi}_{1}(\bar{z},d_{i})$ is the JK sample mean,
@@
We have confirmed that increasing the number of sub-volumes from 50 to 100 does not change the estimated covariance. We conclude that with 50 JK regions, the JK covariance is already well converged and is thus a robust estimate of the measurement covariance.
@@ 3.7 Method to estimate the significance of the detection
Given a set of observations for the dipole and its theoretical prediction, we want to provide a quantitative estimate of its detection. To do this, we determine the detection in units of $\sigma$ as the difference in $\chi^{2}$ between accepting or rejecting the presence of the dipole. In practice, we consider the following $\chi^{2}$ when we include a dipole in our prediction,
@@
We can further consider the $\chi^{2}$ when no dipole is added into the prediction,
@@
It is important to note that the computation of the detection significance requires the inversion of the covariance matrix. However, JK covariances may contain a non-negligible level of noise, which can lead to biased results. We correct for this by applying the Hartlap factor (Hartlap et al. 2007). In more detail, we multiply the inverse covariance matrix by $(N_{\rm JK}-p-2)/(N_{\rm JK}-1)$, where $N_{\rm JK}$ stands for the number of JK regions and $p$ represents the number of separation bins.
@@ 4 Results
Table 2: Specifics of the most representative galaxy samples used in the analysis. Biases have been measured in real space and neglect flux magnification. We report the dimensionless quantity $\overline{\Psi}^{\mathrm{NFW}}_{0}/c^{2}$, where $\overline{\Psi}^{\mathrm{NFW}}_{0}$ corresponds to the mean value of NFW potential at the galaxy positions for the selected objects. We did not consider the split ($50\%$ faint, $50\%$ bright) for the case ‘centrals + satellites’.

### v2

@@ 3.3 Flux magnification
Since the luminosity distance is affected by perturbations (Bonvin et al. 2006; Challinor & Lewis 2011), the measured fluxes are also affected by relativistic effects such as magnification and Doppler corrections. While the observed fluxes in FS include observational effects such as dust extinction, magnification due to lensing and Doppler effects are not included. We corrected the fluxes to account for these effects using the relation
@@
$\mu_{\rm L}=\frac{1}{(1-\kappa)^{2}-\gamma_{1}^{2}-\gamma_{2}^{2}}\,.$ (29)
@@ 3.4 Flux reference to perform split in bright and faint
In order to ensure that bright and faint populations have the same redshift evolution across all bins, we did not use a single flux cut. In fact, flux is inversely proportional to the square of the luminosity distance, meaning that a fixed flux cut would lead to more bright objects at low redshift and more faint objects at high redshift. To avoid this asymmetry in the redshift distribution, we performed the split in $20$ sub-bins, following a similar procedure to the one described in Bonvin et al. (2023).
@@ 3.5 Estimation of biases
We describe here the method we used to estimate the galaxy bias and local count slope for the full Euclid spectroscopic sample, as well as for the bright and faint populations. In our analysis, we neglect the impact of evolution bias [^4] Our galaxy selection leads to the same observed redshift evolution of the two populations and partially cancels out the dipole contribution of source count evolution, which is proportional to the difference in the evolution biases of the two populations. Neglecting evolution bias is also fully justified by the outcome of our analysis in Sect. 4.1, where we show that the overall linear dipole is very small..
@@ 3.5.1 Galaxy bias
The galaxy biases for the full and the bright/faint samples is estimated using the method described in Lepori et al. (2023) and Bonvin et al. (2023). We constructed a map of the galaxy number counts using HEALPix (Górski et al. 2005) and we extracted its angular power spectrum; we fitted it to a theory prediction of the power spectrum computed with the code class (Di Dio et al. 2013, 2014b; Lesgourgues 2011; Blas et al. 2011), using HMCODE to account for nonlinear effects (Mead et al. 2016). The angular power spectrum of the map is estimated with the library NaMaster (Alonso et al. 2019), which corrects for the effect of the mask. The HEALPix map is constructed from the comoving position of the objects, and therefore its angular power spectrum only includes density contributions. Thus, we fitted the galaxy bias in real space, so included only the density term in our theory model. In the fit of the simulated spectrum to the model, we fixed the cosmology so that the only free parameter is the galaxy bias.
@@ 3.5.2 Local count slope
and we introduce the LF of bright and faint galaxies in the redshift bin $\bar{z}$, that is $\bar{N}_{\mathrm{B}}\equiv N_{\mathrm{B}}(\bar{z},\,\bar{L}_{\mathrm{H}\alpha}\geq{L}_{\mathrm{H}\alpha,\,\mathrm{lim}},\,\bar{L}_{Y_{\scriptscriptstyle\rm E}}\geq{L}_{Y_{\scriptscriptstyle\rm E},\,\mathrm{lim}})$ and $\bar{N}_{\mathrm{F}}\equiv N_{\mathrm{F}}(z,\,\bar{L}_{\mathrm{H}\alpha}\geq{L}_{\mathrm{H}\alpha,\,\mathrm{lim}},\,\bar{L}_{Y_{\scriptscriptstyle\rm E}}<{L}_{Y_{\scriptscriptstyle\rm E},\,\mathrm{lim}})$, respectively. The luminosities $\bar{L}_{X}$ are the luminosities corresponding to the flux $F_{X}$ in the background FLRW cosmology.
@@
In practice, we estimated the slopes $s_{\mathrm{B/F},\mathrm{H}\alpha}$ by fixing the flux cuts in the $Y_{\scriptscriptstyle\rm E}$ band and performing the same cuts on a version of the mock that is deeper in $\mathrm{H}\alpha$ flux. The slope $s_{\mathrm{B},Y_{\scriptscriptstyle\rm E}}$ is estimated after applying the selection in the $\mathrm{H}\alpha$ flux, from the cumulative LF of the bright population at the flux cut in the $Y_{\scriptscriptstyle\rm E}$ band. Since we performed the split between bright and faint in sub-bins, the $Y_{\scriptscriptstyle\rm E}$-band flux cut is redshift dependent. We estimated the slope of the cumulative LF in sub-bins and estimated an effective value by taking an average weighted by the number of objects in each sub-bin.
@@ 3.6 Measurements of the dipole and its covariance
We used the Landy–Szalay (LS) estimator (Landy & Szalay 1993) to measure the 3D cross-correlation between bright and faint objects within each redshift bin,
@@
The LS estimator has been shown to have the lowest possible variance and bias among all possible estimators based on pair counts, if the correlation is small ($\xi\ll 1$) and if the number densities of the random catalogues are sufficiently high (Landy & Szalay 1993; Kerscher et al. 2000; Keihänen et al. 2019). In order to get unbiased results, we hence created random catalogues that have ten times as many objects as their corresponding data catalogues.
In practice, we created the Poisson-sampled random catalogues from the density distribution of the data catalogues using the inversion method, as described in Schulz (2023). The LS estimator is computed using a modified version of the publicly available code CUTE (Alonso 2012). The modifications, which are the same as in Breton et al. (2019), extend the capabilities of CUTE to allow the computation of odd multipoles in the two-point correlation function. We counted the pairs for separations $2\,h^{-1}\,\mathrm{Mpc}\leq d\leq 30\,h^{-1}\,\mathrm{Mpc}$ and orientations $-1\leq\mu\leq 1$, binning into $d$-bins of width $\Delta d=2\,h^{-1}\,\mathrm{Mpc}$ and $\mu$-bins of width $\Delta\mu=0.004$.
We did not estimate the correlation for smaller separations because such separations are not well sampled by our catalogues. The average separation between a bright object and the closest faint object can be derived from the cross-correlation function monopole, as it represents the average overdensity of faint objects in the neighbourhood of a bright object. From the measurements of the real-space monopoles on the light cone at small separations we found that there are typically no faint objects separated by $d\lesssim 1\,h^{-1}\,\mathrm{Mpc}$ from the average bright object, and at $\bar{z}=1.0$, the average separation between a bright and the closest faint object is $d=3.6\,h^{-1}\,\mathrm{Mpc}$. The number of pairs with separation below this value is generally small, leading to very large Poisson errors on the measurement at those separations. Additionally, at very small separations, the LS estimator can in principle become undefined in certain ($d,\mu$)-bins if there are no pairs between the two random catalogues. For these reasons, we excluded $d<2\,\,h^{-1}\,\mathrm{Mpc}$ from the analysis.
@@
In practice, the integral is numerically computed as a sum over the $\mu$-bins. We thus computed the dipole of the LS estimator as
@@
The covariance of the measurement was estimated with the jackknife (JK) method (Norberg et al. 2009),
@@
where the survey volume at redshift $\bar{z}$ was divided into $N=50$ sub-volumes with approximately equal surface area and shape by running a kmeans [^5] https://github.com/esheldon/kmeans_radec/ algorithm (Kwan et al. 2017; Suchyta et al. 2016) on one of the corresponding random catalogues. The same split was applied to the data catalogues and the LS dipole was measured $N$ times, each time removing a different sub-volume $k$ from the survey. In Eq. (36), $\xi^{k}_{1}(\bar{z},d_{i})$ is the dipole measured within the survey volume at redshift $\bar{z}$ from which the $k$-th sub-volume has been removed, and $\bar{\xi}_{1}(\bar{z},d_{i})$ is the JK sample mean,
@@
We confirmed that increasing the number of sub-volumes from 50 to 100 does not change the estimated covariance. We conclude that with 50 JK regions, the JK covariance is already well converged and is thus a robust estimate of the measurement covariance.
@@ 3.7 Method to estimate the significance of the detection
Given a set of observations for the dipole and its theoretical prediction, we want to provide a quantitative estimate of its detection. To do this, we determined the detection in units of $\sigma$ as the difference in $\chi^{2}$ between accepting or rejecting the presence of the dipole. In practice, we considered the following $\chi^{2}$ when we include a dipole in our prediction,
@@
We further considered the $\chi^{2}$ when no dipole is added into the prediction,
@@
It is important to note that the computation of the detection significance requires the inversion of the covariance matrix. However, JK covariances may contain a non-negligible level of noise, which can lead to biased results. We corrected for this by applying the Hartlap factor (Hartlap et al. 2007). In more detail, we multiplied the inverse covariance matrix by $(N_{\rm JK}-p-2)/(N_{\rm JK}-1)$, where $N_{\rm JK}$ stands for the number of JK regions and $p$ represents the number of separation bins.
@@ 4 Results
4.1 Relativistic dipole in the linear regime
[…]
4.2 Relativistic dipole in the nonlinear regime
In Fig. 6 we show the measurement of the dipole in the four redshift bins at $\bar{z}=1,1.2,1.4,1.65$, for the $50\%$-$50\%$ case (left panels) and $10\%$-$90\%$ case (right panels). The catalogue processed to obtain these measurements only contains central galaxies, and we neglect flux magnification. The predictions of theory is estimated from the model in Sect. 2.2, which assumes that the kinematic contributions to the dipole are adequately modelled using linear theory, with nonlinearities mainly affecting the gravitational redshift contribution. The specifics of the galaxy populations (galaxy bias and mean gravitational potential at the galaxy position) are estimated from the catalogues and their values are reported in Table 4.
Table 2: Detection significance.
theory covariance JK covariance combined covariance
Redshift $50\%$ $50\%$ $10\%$ $90\%$ $50\%$ $50\%$ $10\%$ $90\%$ $50\%$ $50\%$ $10\%$ $90\%$
$1$ $6.5\sigma$ $9.2\sigma$ $3\sigma$ $4\sigma$ $3.1\sigma$ $4.5\sigma$
$1.2$ $5.8\sigma$ $7.3\sigma$ $2.9\sigma$ $3.4\sigma$ $3.2\sigma$ $3.7\sigma$
$1.4$ $5\sigma$ $1.6\sigma$ $2.3\sigma$ No detection $2.5\sigma$ $1.5\sigma$
$1.65$ $4.9\sigma$ $6\sigma$ $2.2\sigma$ $2.1\sigma$ $2.2\sigma$ $2.1\sigma$
666 We report the detection significance in each redshift bin, for the $50\%$-$50\%$ split and the $10\%$-$90\%$ split. The columns labelled with ‘theory covariance’ use the linear covariance described in Sect. 2.3. The columns labelled with ‘JK covariance’ use the numerical covariance estimated with the JK method, as detailed in Sect. 3.6. The ‘combined covariance’ corresponds to the rescaling of the theory covariance matrix by the diagonal elements of the JK covariance.
where ${\rm Corr}^{\rm th}_{ij}$ is the theory correlation matrix, given by
and $i$ and $j$ run over the components of the covariance matrix. The combined covariance provides the same error estimate as the JK method, having the advantage that the non-diagonal elements do not exhibit spurious numerical oscillations and follow the trend expected from the theory covariance. Assuming the measurements in different redshift bins are independent, we also computed the total detection significance by summing in quadrature the significance of all bins. The first data point at $d=1\,h^{-1}\,\mathrm{Mpc}$ is excluded from this analysis.
The estimation of the detection significance assumes perfect knowledge of the linear galaxy bias of the two populations. While we can measure this quantity very accurately in simulations, where the underlying cosmological model is known, this will not be possible with the Euclid real data. This issue can be overcome by considering a joint analysis of the even multipoles of both galaxy populations combined with the dipole measurement, leaving the galaxy bias as free parameters in the analysis. It is shown in Sobral-Blanco & Bonvin (2022) that the bias of the bright and faint populations are typically very well constrained by the even multipoles and that the uncertainty on the biases barely impacts the measurement of gravitational redshift. More precisely, comparing the detectability of gravitational redshift with the detectability of gravitational redshift multiplied by the bias difference $b_{\rm B}-b_{\rm F}$ leads to a difference in precision of 0.1%. This shows that even if the bias difference is fully degenerated with the gravitational redshift signal in the dipole, the biases are so well constrained by the even multipoles that the degeneracy can be extremely well broken.
4.3 Isolating the different contributions to the dipole
In Fig. 7 we show the measurements for a) and b) in the left panels, and for c) and d) in the right panels, compared to our theory prediction that includes linear velocities and nonlinear gravitational potential. The two contributions of peculiar velocities, coming from the light-cone and Doppler effects, are individually consistent with zero at $\bar{z}=1.2,1.4,1.65$. In the redshift bin $\bar{z}=1$, the light-cone effect is not completely negligible, as the dipole measurements exhibit a systematic negative amplitude on scales below $10\,h^{-1}\,\mathrm{Mpc}$. However, the combined Doppler and light-cone effects lead to a dipole consistent with zero on scales $>3\,h^{-1}\,\mathrm{Mpc}$. This suggests that at $\bar{z}=1$ the contributions from nonlinear Doppler effects and the light-cone effects are individually not negligible. However, since they generate a dipole of opposite sign, their combined contribution is consistent with zero on scales $>3\,h^{-1}\,\mathrm{Mpc}$. Since both of these effects are coupled to the gravitational redshift in the full dipole measurements, we cannot robustly conclude that the effect of nonlinear peculiar velocities is negligible at $\bar{z}=1$, where we forecast the greater detection significance. A model that can consistently treat both nonlinear velocities and nonlinear gravitational potential is needed for a correct interpretation of the dipole measurement at these scales.
4.4 Impact of satellite galaxies and flux magnification
In our baseline settings, we have only included central galaxies, and we have neglected flux magnification. In this section, we discuss the impact of satellite galaxies and flux magnification on the dipole measurements, which includes both the Doppler and the subdominant contribution from lensing magnification.
We performed the dipole measurements for our baseline split ($10\%$-$90\%$) including satellite objects in the sample (and neglecting flux magnification). The specifics for the catalogues with satellites are reported in Table 4. Compared to the selection with only central objects, including satellites results in a larger number of objects, larger galaxy bias, and higher average gravitational potential at the positions of the sources for both populations. This is because satellite galaxies are primarily found inside the most massive haloes of the simulation.
Table 3: Impact of satellite galaxies and flux magnification on the detection significance.
$1$ $4.5\,\sigma$ $4.4\,\sigma$ $4.7\,\sigma$
$1.2$ $3.7\,\sigma$ $2.8\,\sigma$ $3.4\,\sigma$
$1.4$ $1.5\,\sigma$ $<1\,\sigma$ $1.4\,\sigma$
$1.65$ $2.1\,\sigma$ $1.8\,\sigma$ $1.8\,\sigma$
777 For this estimate, we have considered the $10\%$ bright - $90\%$ faint split and used the ‘combined’ covariance.
We perform a similar analysis to test the impact of flux magnification. In the right panels of Fig. 8 we compare the dipole measurements for the cases with and without flux magnification. Flux magnification does not affect the gravitational redshift contribution to the dipole, and since this is by far the largest contribution to our prediction, its impact on the theory model is largely negligible. The measured dipole for the catalogues with flux magnification is consistent with the measurements without flux magnification. In Table 3 we report the significance of the detection for these measurements. We find that it is compatible with the estimate for our baseline setting, with only central galaxies and no flux magnification. The differences between the two cases are $\leq 0.3\,\sigma$ in all redshift bins.
5 Conclusions
In this paper, we present a measurement of the dipole of the two-point correlation function in the Flagship simulation, for a selection of objects tailored to the Euclid Spectroscopic Survey. This observable is sourced by gravitational redshift and Doppler effects, and it has been shown that it can be used to test the equivalence principle on cosmological scales, see, for example, Bonvin & Fleury (2018).
Since the contribution of gravitational redshift to the observed redshift is not currently included in the Flagship mock, we implemented this effect by modelling the gravitational potential at the position of each galaxy using the NFW profile for the host haloes. In order to isolate these relativistic effects from the standard clustering contributions, given by the galaxy overdensity and RSDs, we split the full sample of galaxies into two populations. The split is performed using their fluxes in the $Y_{\scriptscriptstyle\rm E}$ band as reference. We show that this choice leads to larger differences in galaxy bias for the bright and faint sample, being more correlated with the mass of the host halo than the ${\rm H}\alpha$ flux. On large scales, the dipole measured from the Flagship mock is consistent with zero regardless of the percentage of bright and faint objects. This is not surprising, as a signal-to-noise analysis based on linear theory leads to $\text{S}/\text{N}<1$. On small scales, the amplitude of the signal is significantly larger due to the contributions of nonlinear gravitational redshift. Using the model presented in Saga et al. (2020, 2022) as a reference, on scales $d<30\,h^{-1}\,{\rm Mpc}$ and for our fiducial split with $90\%$ faint objects and $10\%$ bright objects, we found a detection significance of $4.5\,\sigma$, $3.7\,\sigma$, $1.5\,\sigma$, and $2.1\,\sigma$ in the four redshift bins $\bar{z}=1,1.2,1.4,1.65$, respectively. The overall detection significance amounts to $6\,\sigma$. These values are obtained from the combined covariance, which rescales the analytical covariance to match the uncertainty estimated from the mock using the JK method. This is our fiducial choice, which leads to a conservative estimate. A split with equal number of bright and faint objects leads to a slightly lower detection significance at low redshift.
While our nonlinear model accounts for nonlinear gravitational redshift, it models velocities using linear theory. We tested this assumption by separating the different contributions to the dipole: the light-cone effect, the sum of light-cone and Doppler effects, and the gravitational redshift. Although the gravitational redshift is the largest contribution to the nonlinear dipole, the light-cone effect is not completely negligible at $\bar{z}=1$. Nevertheless, the contributions from light-cone and Doppler effects appear to roughly cancel out. We cannot interpret this cancellation in our nonlinear model, since it only models velocities using linear theory. To correctly separate the different contributions to the dipole at low redshift and on small scales, we require a model that can consistently treat both the nonlinearities in the velocities and the gravitational potential (see, e.g., Dam & Bonvin 2023). We leave the development of a complete nonlinear model to future work.
We also studied the impact of satellite galaxies and flux magnification on the estimated dipole, finding that the detection significance of the dipole is not substantially affected by them. Although flux magnification has negligible impact on our measurements, in Appendix A we validate different methods to estimate the local count slope for selections with multiple flux cuts. This work is relevant for modelling lensing magnification in galaxy clustering when the selection of galaxies is not a simple flux cut. The measurements in the Flagship mock catalogue cover $1/8$ of the sky, which roughly corresponds to $5000\,{\rm deg}^{2}$. This is about twice the sky coverage planned for Euclid DR1. Assuming that the number density and galaxy bias of the Flagship mock reflects the properties of the galaxies observed by Euclid in all the three planned Data Releases (Euclid Collaboration: Scaramella et al. 2022; Euclid Collaboration: Mellier et al. 2024), we forecast that a detection of the dipole is foreseeable for DR2 and DR3. In estimating the detection significance of the dipole, we have assumed that the Euclid spectroscopic sample is pure and complete. A detailed analysis accounting for these effects and other systematic uncertainties is left for future work.
While we include in our measurements scales $d<30\,h^{-1}\,{\rm Mpc}$, we point out that most of the signal-to-noise comes from separations $d\sim 5\,h^{-1}\,{\rm Mpc}$. This is the typical scale at which gravitational redshift has been detected in galaxy clusters; see, for example, Wojtak et al. (2011)Sadeh et al. (2015), and Rosselli et al. (2023). The key observable in these measurements is the distribution of line-of-sight (LOS) velocities for cluster member galaxies around the cluster centre: the kinematic Doppler effect gives rise to a broad symmetric distribution around a zero LOS velocity, while gravitational redshift shifts the value of the mean velocity of this distribution. Compared to the cluster measurements, the method we propose has the advantage of not requiring the identification of cluster member galaxies or associating the cluster centre with a specific galaxy in the spectroscopic catalogue. Measuring the galaxy cross-correlation function is, in a way, more general, as various splitting strategies can be tested and combined to optimise the measurements. The different systematics affecting the two observables make them highly complementary. A quantitative comparison between the two approaches would require forecasting the detectability of gravitational redshift in the Euclid cluster catalogue (Sartoris et al. 2016; Euclid Collaboration: Adam et al. 2019), which is beyond the scope of this work.
Acknowledgements.
The work of FL, SS, JA, and CB is supported by the Swiss National Science Foundation. CB and LD acknowledge support from the European Research Council (ERC) under the European Union’s Horizon 2020 research and innovation program (grant agreement No. 863929; project title “Testing the law of gravity with novel large-scale structure observables”). The Euclid Consortium acknowledges the European Space Agency and a number of agencies and institutes that have supported the development of Euclid, in particular the Agenzia Spaziale Italiana, the Austrian Forschungsförderungsgesellschaft funded through BMK, the Belgian Science Policy, the Canadian Euclid Consortium, the Deutsches Zentrum für Luft- und Raumfahrt, the DTU Space and the Niels Bohr Institute in Denmark, the French Centre National d’Etudes Spatiales, the Fundação para a Ciência e a Tecnologia, the Hungarian Academy of Sciences, the Ministerio de Ciencia, Innovación y Universidades, the National Aeronautics and Space Administration, the National Astronomical Observatory of Japan, the Netherlandse Onderzoekschool Voor Astronomie, the Norwegian Space Agency, the Research Council of Finland, the Romanian Space Agency, the State Secretariat for Education, Research, and Innovation (SERI) at the Swiss Space Office (SSO), and the United Kingdom Space Agency. A complete and detailed list is available on the Euclid web site (www.euclid-ec.org). This work has made use of CosmoHub (Tallada et al. 2020; Carretero et al. 2017). CosmoHub has been developed by the Port d’Informacio Cientifica (PIC), maintained through a collaboration of the Institut de Fisica d’Altes Energies (IFAE) and the Centro de Investigaciones Energeticas, Medioambientales y Tecnologicas (CIEMAT) and the Institute of Space Sciences (CSIC & IEEC). CosmoHub was partially funded by the ”Plan Estatal de Investigacion Cientifica y Tecnica y de Innovacion” program of the Spanish government, has been supported by the call for grants for Scientific and Technical Equipment 2021 of the State Program for Knowledge Generation and Scientific and Technological Strengthening of the R+D+i System, financed by MCIN/AEI/ 10.13039/501100011033 and the EU NextGeneration/PRTR (Hadoop Cluster for the comprehensive management of massive scientific data, reference EQC2021-007479-P) and by MICIIN with funding from European Union NextGenerationEU(PRTR-C17.I1) and by Generalitat de Catalunya.
Appendix A Validation of the effective local count slope estimates
Using that
The parameters $s_{\mathrm{B},\mathrm{H}\alpha}$, and $s_{\mathrm{B},Y_{\scriptscriptstyle\rm E}}$ are defined as
where
$\left.\frac{\partial\bar{N}_{\mathrm{F}}}{\partial\ln\bar{L}_{Y_{\scriptscriptstyle\rm E}}}\right\rvert_{{L}_{Y_{\scriptscriptstyle\rm E},\,\mathrm{lim}}}=-\left.\frac{\partial\bar{N}_{\mathrm{B}}}{\partial\ln\bar{L}_{Y_{\scriptscriptstyle\rm E}}}\right\rvert_{{L}_{Y_{\scriptscriptstyle\rm E},\,\mathrm{lim}}}\,,$ (51)
and thus
[…]
  1. 1.
    $F_{\mathrm{H}\alpha}\rightarrow\tilde{F}_{\mathrm{H}\alpha}=\tilde{\mu}_{\rm L}\,F_{\mathrm{H}\alpha},\qquad F_{Y_{\scriptscriptstyle\rm E}}\rightarrow\tilde{F}_{Y_{\scriptscriptstyle\rm E}}=\tilde{\mu}_{\rm L}\,F_{Y_{\scriptscriptstyle\rm E}}\,.$ (55)
  2. 2.
  3. 3.
We compared the value of the local count slope using the two methods described above, finding agreement at the level of a few percent. We tested the consistency of both these measurements with the lensing magnification measured in the cross-correlation between redshift bins. As test cases, we considered the split with $50\%$ bright and $50\%$ faint objects (split 1), and the split with $10\%$ bright and $90\%$ faint objects (split 2). We constructed number count maps of the bright and faint populations. The binning in redshift was performed using the background redshift, so that RSD effects are not present. We considered two types of maps: maps that are constructed from the comoving position of the objects and selected based on their intrinsic flux (no flux magnificaton), and maps constructed from the magnified position of the objects and selected based on their magnified flux. The first case includes only contribution from the local density fluctuations of the galaxies, and the second case includes also lensing magnification.
For each split, we estimated the angular cross spectrum by cross-correlating a map at low redshift ($\bar{z}=1$) with a map at high redshift ($\bar{z}=1.65$). In total, there are four possible configurations: (map of bright galaxies at $\bar{z}=1$) $\times$ (map of bright galaxies at $\bar{z}=1.65$), (map of faint galaxies at $\bar{z}=1$) $\times$ (map of bright galaxies at $\bar{z}=1.65$), (map of bright galaxies at $\bar{z}=1$) $\times$ (map of faint galaxies at $\bar{z}=1.65$), and (map of faint galaxies at $\bar{z}=1$) $\times$ (map of faint galaxies at $\bar{z}=1.65$). Following Fosalba et al. (2015), we constructed an estimator for the magnified spectra $C^{\mathrm{est}}_{\ell}$, given by
Appendix B Properties of the galaxy populations
In Table 4, we report the properties of the galaxy populations we have used in our analysis.
Table 4: Specifics of the most representative galaxy samples.